# ShareChat: A Dataset of Chatbot Conversations in the Wild

### v1

@@ 3.4 Topic Analysis
We classify each user message into one of 24 predefined topic categories using multilingual LLMs Llama-3.1-8B-Instruct 444https://huggingface.co/meta-llama/Llama-3.1-8B under a few-shot prompting setup. To enable large-scale processing, we run inference with 4-bit quantization, batch size 32. Model outputs are normalized through a post-processing pipeline to handle formatting variants and stray meta-responses, ensuring reliable mapping to our fixed taxonomy. For downstream analyses, the 24 fine-grained categories are further consolidated into seven higher-level groups to support both granular and coarse-grained comparisons across platforms. The 7 high-level categories of user requests are plot in the Figure. 4.
Overall, the averaged distribution reveals that information seeking remains the core driver of LLM interactions, with creative, technical, and advisory uses forming important but secondary dimensions of user behavior. Specifically, we observe a strong central tendency in how users engage with large language models. Seeking Information overwhelmingly dominates usage, comprising 39.6% of all requests, underscoring that fact-oriented question-answering remains the primary interaction paradigm. The second-largest category, Other/Unknown (19.0%), reflects a substantial proportion of prompts that are either ambiguous, exploratory, or insufficiently specified for reliable categorization. Technical Help (12.4%) and Writing (10.1%) constitute the next major cluster of user intent, highlighting the continued reliance on LLMs for procedural support and text generation tasks. Practical Guidance (8.4%) and Self-Expression (8.1%) appear at comparable frequencies, indicating moderate but consistent engagement with step-by-step advice and personal or expressive communication. Finally, Multimedia accounts for only 2.4% of requests, likely reflecting limited multimodal capabilities at the time of data collection. Detailed prompts, category definitions, and classification statistics are provided in the Appendix A.3.
4 Use Cases
@@ 4.1 Conversation Completeness Analysis
To assess how comprehensively large language models fulfill user needs in multi-turn help-seeking dialogues, we developed an automated completeness analysis pipeline that evaluates whether each user intention across a conversation is adequately addressed by the assistant. Our analysis reveals distinct performance across platforms, with ChatGPT and Claude demonstrating the highest consistency in fully satisfying user intentions, while Gemini, Grok and Perplexity exhibit significantly wider variance in conversational quality across multi-turn interactions.
The pipeline implements a three-stage workflow adapted from the DeepEval555https://github.com/confident-ai/deepeval?tab=readme-ov-file conversation completeness metric framework. First, we extract user intentions from each conversation by prompting Qwen3-8B Yang et al. (2025) with a temperature of 0.7 to identify the distinct goals or information needs expressed in the user turns, yielding a chronologically ordered list of intentions per conversation. Second, for each identified intention, we construct an evaluation prompt that presents the full conversation context alongside the specific intention and instruct Qwen3-8B to classify whether that intention was satisfied; the model produces a categorical verdict of complete (the assistant fully addressed the request), partial (the assistant provided some relevant information but the goal was incompletely met), or incomplete (the assistant failed to address the request or the conversation ended before meaningful progress). Finally, we aggregate these intention-level verdicts into a conversation-level completeness score by computing the weighted proportion of intentions, where complete verdicts contribute 1.0, partial verdicts contribute 0.5, and incomplete verdicts contribute 0.0. This three-way categorical framework allows us to capture nuanced degrees of conversational success beyond binary outcomes. We apply this pipeline uniformly across all five platforms in ShareChat, enabling direct cross-platform comparisons of how effectively different assistants satisfy user needs over the course of extended interactions. The prompts used and experiment details could be found in Appendix A.4.
@@
Different platforms have revealed substantial variation in how effectively they address user intentions in multi-turn conversations. ChatGPT and Claude exhibit the highest rates of complete verdicts in Figure 5(a), with both platforms resolving the majority of user intentions fully. Gemini and Grok show more balanced distributions between complete and partial verdicts, with a substantial proportion of intentions receiving incomplete verdicts. Perplexity demonstrates an intermediate pattern with the highest rate of partial verdicts across all platforms. These intention-level patterns are also reflected in the conversation-level completeness scores (Figure 5(b)), where ChatGPT, Claude, Grok, and Perplexity all achieve median scores of 1.0, while Gemini’s median is 0.833.
Also, the number of extracted intentions per conversation (Figure 5(c)) varies systematically across platforms, with ChatGPT and Claude conversations containing a median of 2 intentions, while Gemini, Grok, and Perplexity show a median of 1 intention per conversation, which aligns with the results in Figure 8.
These findings demonstrate the structural diversity of interactions in the dataset, ranging from single-turn queries to complex sessions involving multiple distinct intentions. Furthermore, the prevalence of partial verdicts highlights that many real-world interactions result in incomplete goal satisfaction, a dynamic often obscured in benchmarks that prioritize single-turn correctness over multi-turn resolution. It showcases the functional emphasis of each platform, as the higher intention counts and completeness scores in ChatGPT and Claude reflect their usage for complex or general purpose problem solving, while the frequent partial verdicts and lower intention depth in Perplexity correspond to its search oriented design where model outputs often serve as intermediate information retrieval rather than exhaustive answers. By providing completeness labels alongside the raw conversation data, ShareChat supports the development of automatic evaluation metrics that better align with user satisfaction in multi-turn interactions, which in turn can serve as valuable training signals for reinforcement learning.
4.2 Response Source Analysis
From the statistics, the Grok dataset contains 14,415 conversations, with 8,242 (57.18%) including sources and 6,173 (42.82%) without sources. The Perplexity dataset includes 17,305 conversations, with 8,545 (49.38%) containing sources and 8,760 (50.62%) without sources. More detailed information of these two platforms are provided in Tabel 3.
Plot 6(a) shows that Grok typically uses very few sources per conversation, whereas Perplexity exhibits a long-tailed distribution in which many conversations draw on dozens of sources. In addition, Grok treats “X” as a special source, likely due to privileged access to that platform. This could yield more timely but also potentially noisier or less reliable information, suggesting that the use of X as a primary source warrants further examination for possible misinformation risks. Plot 6(b) further demonstrates Grok’s heavy dependence on this single dominant source (“X”), producing a highly skewed distribution. In contrast, Perplexity draws from a much broader set of sources, with English Wikipedia being the most frequent but not overwhelmingly dominant.
@@ 4.2 Response Source Analysis
Taken together, these plots demonstrate that the two platforms employ fundamentally distinct information retrieval strategies, prioritizing different types of external evidence to groud their generated responses: Grok gravitates toward X, which is more social-media-driven and volatile, while Perplexity leans toward Wikipedia, which is more curated, structured, and stable.
@@ 4.3 Timestamp Analysis
Timestamp data for the Grok and ChatGPT datasets permits a comprehensive temporal analysis as shown in Figure 7. Timestamps were present for the vast majority of turns (99.97% for ChatGPT, 100% for Grok). For the small fraction of missing timestamps in ChatGPT data (1.40% of turns), we employed linear interpolation between known timestamps where feasible, and dropped affected turns when interpolation was not possible. We verified that results remained stable across different handling strategies, confirming that the small amount of missing data did not bias our findings.
After removing extreme outliers above the 99th percentile, we observe systematic differences between platforms. ChatGPT exhibits longer mean user response times than Grok (1,580 vs. 931 seconds), while producing shorter mean LLM response times (18.4 vs. 24.6 seconds). Median values are substantially closer across platforms (users: 93.4 vs. 91.2 seconds; LLMs: 6.9 vs. 16.8 seconds), indicating that mean differences are driven primarily by long-tail behavior.
To assess whether longer LLM outputs demand increased user processing time, we analyze the relationship between LLM response length (in characters) and subsequent user response latency. Across both platforms, correlations are near zero (ChatGPT: ; Grok: ), indicating that verbosity provides little predictive signal for typical user response behavior. While longer responses permit longer delays, they do not systematically increase central tendency, challenging implicit verbosity-as-engagement assumptions. The details could be found in Appendix A.5.
Finally, we examine how response times evolve over the course of a conversation by correlating turn position with response latency. ChatGPT exhibits a significant negative correlation between turn index and LLM response time (Pearson ; Spearman ), indicating progressively faster responses as conversations lengthen, consistent with caching, adaptation, or context optimization effects. In contrast, Grok shows a significant positive correlation (Pearson ; Spearman ), suggesting increasing computational overhead with conversation length. User response times show minimal linear association with turn position on both platforms (ChatGPT: , ; Grok: , ). Together, these divergent temporal dynamics highlight fundamental architectural differences in how the two systems handle extended interactions.
@@ 5 Discussion
The findings presented in this study offer several important implications for understanding human-AI interaction and for guiding future research on conversational LLMs. We structure our discussion around four key dimensions enabled by ShareChat. While we only have demonstrated three concrete use cases: conversation completeness, source content analysis and temporal analysis, the dataset’s utility extends well beyond these applications.
@@
Complex Interaction Dynamics in Extended Contexts. ShareChat’s longer conversation contexts, averaging 4.62 turns compared to 2.02 in LMSYS-Chat-1M Zheng et al. (2024), provide a critical resource for studying phenomena that emerge only over extended interactions. Prior work has shown that model reliability can degrade as instructions are refined across turns Laban et al. (2025b), yet most existing corpora offer limited opportunities to investigate such dynamics due to their predominantly single-turn structure. The presence of conversations extending into dozens of turns, combined with the diversity of intention counts per conversation, enables researchers to examine how user needs evolve, how models maintain coherence over long contexts, and when breakdowns occur.
ShareChat makes it possible to study these partial successes systematically, revealing where assistants provide useful but insufficient information and how users adapt their follow-up requests in response. This diversity can inform the development of models better equipped to manage multiple user needs within a single interaction, learning when to shift focus, when to return to unresolved intentions, and how to maintain coherence while serving users whose goals evolve over time.
@@
Platform-Dependent Usage and Performance Variability. Unlike the “universal assistant” assumption often found in general benchmarks, our cross-platform analysis reveals that user behavior is highly sensitive to the specific design and functional positioning of each platform. Our topic distribution analysis demonstrates that users actively select platforms for distinct roles, which aligns their requests with perceived strengths, such as using Perplexity for information retrieval versus Claude for technical reasoning. This functional specialization is mirrored in conversation completeness: while ChatGPT and Claude exhibit narrow, concentrated completeness scores indicating highly predictable output, Gemini and Grok show substantial variability. Notably, the high frequency of “partial” verdicts in Perplexity interactions reflects its distinct role as a search engine, where outputs often serve as intermediate steps rather than final answers. Researchers can use ShareChat to investigate these specialized performance profiles, moving beyond generic comparisons to determine which architectural features drive reliability for specific user intents.
@@
Evidence-Based Reasoning and Temporal Metadata. The preservation of rich metadata enables future research on retrieval-augmented generation (RAG) and interaction pacing in naturalistic settings. Our source content analysis reveals pronounced differences in how platforms integrate external evidence. Grok relies overwhelmingly on concentrated source types, while Perplexity draws from a diversified set of domains. These patterns have implications for understanding how users interpret and rely on cited evidence, and how platform affordances shape trust and verification behaviors. Additionally, the availability of turn-level timestamps enables the investigation of conversational rhythm and interaction pacing, offering a temporal dimension of user experience that text-only corpora cannot provide. Our analysis exposes contrasting architectural dynamics: while ChatGPT demonstrates progressively lower latency as conversations extend over multiple turns, Grok exhibits an inverse trend, with users experiencing slower generation speeds as context length grows.
@@
Self-Selection and Safety in Public Sharing. Finally, the substantially lower toxicity rates in our corpus compared to existing benchmarks, for both user messages and LLM responses, likely reflect the self-selection inherent in public sharing. Users who choose to share conversations may filter for interactions they consider appropriate or valuable, producing a corpus that captures more constructive usage patterns. This suggests that while public datasets may not fully capture the “dark side” of LLM usage (such as jailbreaking attempts), they offer a cleaner, high-quality resource for studying successful task completion and constructive human-AI collaboration.
@@ 6 Conclusion
In this study, we addressed the limitations of existing single interface datasets by constructing ShareChat, a large scale, multi platform corpus of user LLM conversations. By compiling more than 140,000 publicly shared interactions from five distinct platforms, we provided a resource that preserves the native metadata and diverse interaction designs of the current AI ecosystem. Our examination of the dataset statistics shows that the self selection process inherent in public sharing leads to a corpus with significantly longer conversation contexts and lower toxicity compared to standard interaction logs.
We further illustrated the value of this data through three specific analysis. First, our conversation completeness analysis indicated that user intentions are more frequently satisfied in general purpose models like ChatGPT and Claude, reflecting their emphasis on broad capabilities. Second, our source content analysis uncovers the distinct information retrieval strategies employed by Grok and Perplexity. Third, our temporal analysis suggests that ChatGPT and Grok exhibit contrasting temporal dynamics. Collectively, these findings highlight pronounced differences in user behavior that align with the functional emphasis of each platform, underscoring the importance of studying LLM capabilities within their specific ecological contexts.
ShareChat offers broader utility for connecting model capabilities with authentic usage patterns. It not only provides a testbed for evaluating model reliability across turns, but also offers rich supervision for training models to handle complex, multi-turn user needs in different settings.
@@
Selection Bias. First, the reliance on publicly shared URLs introduces a trade-off between two forms of bias. On one hand, this approach mitigates the Observer Bias which we talked about in Section 1, where participants alter their behavior simply because they know they are being monitored McCambridge et al. (2014). On the other hand, it introduces Self-Selection Bias Heckman (1979). Because users voluntarily choose which conversations to generate links for, the corpus is likely skewed towards interactions they deem successful, interesting, or novel. Consequently, ShareChat may underrepresent mundane utility tasks or failed interactions where the model performed poorly, as users are less motivated to share such content publicly Hu et al. (2017).
@@
Proxy Nature of Automated Metrics. Finally, while our completeness analysis utilizes robust LLM-based evaluation, it remains a proxy metric. Automated judges provide a scalable approximation of user satisfaction but cannot perfectly capture the nuanced, subjective quality of human preference. These results should be interpreted as strong signals of structural completeness rather than absolute ground-truth measurements of user happiness. Future research is needed to validate these automated metrics against human judgment and to fully explore the evolving longitudinal dynamics of user behavior.
Preliminary Scope of Analysis. Third, the specific analyses presented in this work, such as the source domain distribution and temporal engagement patterns, are intended as illustrative use cases to demonstrate the dataset’s utility. They are not exhaustive sociological studies. Future research is needed to fully explore these dimensions, particularly how these patterns evolve longitudinally as platform features change.
@@
Privacy Preservation and De-Identification. Despite the public provenance of the data, we prioritized the protection of user privacy through a rigorous de-identification pipeline designed to mitigate risks associated with inadvertently exposed sensitive information.
PII Redaction: We utilized automated Named Entity Recognition (NER) and regular expression matching to detect and redact Personally Identifiable Information (PII)—including real names, email addresses, phone numbers, and IP addresses—for different languages.
User Identifier Anonymization: We do not store or release original platform-specific user IDs or usernames. All user identifiers were cryptographically hashed and converted into randomized, internal project identifiers during the initial processing stage. This ensures that the published dataset cannot be easily linked back to user profiles on the original platforms.
Aggregated Analysis: The analyses presented in this paper, like the toxicity assessments and completeness scoring, are conducted solely on aggregated statistics. No individual user’s conversation history is singled out for qualitative scrutiny in a manner that could lead to re-identification or personal embarrassment.

### v2

@@ 3.3 Topic Analysis
We classify each user message into one of 24 predefined topic categories using multilingual LLMs Llama-3.1-8B-Instruct 333https://huggingface.co/meta-llama/Llama-3.1-8B under a few-shot prompting setup. To enable large-scale processing, we run inference with 4-bit quantization, batch size of 32. Model outputs are normalized through a post-processing pipeline to handle formatting variants and stray meta-responses, ensuring reliable mapping to our fixed taxonomy. For downstream analyses, the 24 fine-grained categories are further consolidated into seven higher-level groups to support both granular and coarse-grained comparisons across platforms. The 7 high-level categories of user requests are plotted in Figure 4.
Overall, the averaged distribution reveals that information seeking remains the core driver of LLM interactions, with creative, technical, and advisory uses forming important but secondary dimensions of user behavior. Specifically, we observe a strong central tendency in how users engage with LLMs. Seeking Information overwhelmingly dominates usage, comprising 39.6% of all requests, underscoring that fact-oriented question-answering remains the primary interaction paradigm. The second-largest category, Other/Unknown (19.0%), reflects a substantial proportion of prompts that are either ambiguous, exploratory, or insufficiently specified for reliable categorization. Technical Help (12.4%) and Writing (10.1%) constitute the next major cluster of user intent, highlighting the continued reliance on LLMs for procedural support and text generation tasks. Practical Guidance (8.4%) and Self-Expression (8.1%) appear at comparable frequencies, indicating moderate but consistent engagement with step-by-step advice and personal or expressive communication. Finally, Multimedia accounts for only 2.4% of requests, likely reflecting limited multimodal capabilities at the time of data collection.
Also, we find that users employ platforms for distinct roles, such as using Perplexity for “seeking information” versus Claude for “technical thinking”. Detailed prompts, category definitions, classification statistics and platform-specific results are provided in the Appendix A.5.
4 Representative Analyses
@@ 4.1 Conversation Completeness Analysis
To assess how comprehensively LLMs fulfill user needs in multi-turn dialogues, we develop an automated completeness analysis pipeline that evaluates whether each user intention across a conversation is adequately addressed by the assistant. Our analysis reveals distinct performance across platforms, with ChatGPT and Claude demonstrating the highest consistency in fully satisfying user intentions, while Gemini, Grok and Perplexity exhibit significantly wider variance in conversational quality across multi-turn interactions.
To assess the conversation completeness, we implement a three-stage pipeline using Qwen3-8B yang2025qwen3. First, we extract distinct user intentions from the conversation history. Second, we classify the resolution of each intention as complete, partial, or incomplete. Finally, we compute a conversation-level completeness score by weighting these verdicts. We sample 100 conversations (20 per platform) for human evaluation. Our coding revealed that 85.1% of user intentions are correctly identified, and among these, 92.5% are correctly labeled. Through this annotation scheme, the number of extracted intentions per conversation varies systematically across platforms, with ChatGPT and Claude conversations containing a median of 2 intentions, while Gemini, Grok, and Perplexity show a median of 1 intention per conversation. Full prompts and implementation details are provided in Appendix A.6.
@@
Different platforms reveal substantial variation in how effectively they address user intentions in multi-turn conversations. ChatGPT and Claude exhibit the highest rates of complete verdicts in Figure 5, with both platforms fully resolving the majority of user intentions. Gemini and Grok show more balanced distributions between complete and partial verdicts, with a substantial proportion of intentions receiving incomplete verdicts. Perplexity demonstrates an intermediate pattern with the highest rate of partial verdicts across all platforms.
These findings highlight the dataset’s structural diversity, spanning single-turn queries to complex, multi-intention sessions. The prevalence of partial verdicts reveals that real-world goals often remain incompletely satisfied, which is a dynamic obscured by benchmarks prioritizing single-turn correctness. By providing these completeness labels, ShareChat enables the development of metrics aligned with user satisfaction and offers rich signals for reinforcement learning.
4.2 Source Grounding Analysis
We observe distinct information retrieval paradigms between the two search-enabled platforms. As detailed in Table 2, external grounding is a core feature for both systems: the Grok dataset contains 14,415 conversations with 8,242 (57.18%) including sources, while the Perplexity dataset includes 17,305 conversations with 8,545 (49.38%) containing sources.
As illustrated in Figure 6, Grok exhibits a hyper-concentrated reliance on social data. Its source distribution is heavily skewed toward the X platform, which accounts for 40,624 references, nearly triple the citations for Wikipedia (12,507). This suggests a design optimized for real-time, social-media-driven context, potentially offering higher timeliness but introducing risks related to information stability and bias.
In contrast, Perplexity demonstrates a diversified, encyclopedic retrieval strategy. Its distribution is significantly less concentrated, with English Wikipedia as the leading source (2,919 citations), followed by community forums like Reddit (1,642) and authoritative scientific repositories like NIH (1,339). Furthermore, as detailed in Appendix A.7, Perplexity supports much deeper research sessions, with some conversations integrating over 1,000 distinct sources, whereas Grok typically cites fewer than 20 sources per interaction.
@@ 4.2 Source Grounding Analysis
Taken together, these plots demonstrate that the two platforms employ fundamentally distinct information retrieval strategies, prioritizing different types of external evidence to ground their generated responses.
@@ 4.3 Timestamp Analysis
Timestamp metadata is available for ChatGPT and Grok, covering nearly all turns (99.97% for ChatGPT and 100% for Grok). For the small fraction of missing timestamps in ChatGPT data, we employ linear interpolation between known points where feasible, and dropped affected turns when interpolation was not possible. We verify that results remained stable across different handling strategies, confirming that the missing data does not bias our findings. After removing extreme outliers above the 99th percentile, we observe systematic differences between platforms. ChatGPT exhibits longer mean user response times than Grok (1,580 vs. 931 seconds), while producing shorter mean LLM response times (18.4 vs. 24.6 seconds).
Then, we examine how response times evolve over the course of a conversation by correlating turn position with response latency. ChatGPT exhibits a significant negative correlation between turn index and LLM response time (Pearson ; Spearman ), indicating progressively faster responses as conversations lengthen, consistent with caching, adaptation, or context optimization effects. In contrast, Grok shows a significant positive correlation (Pearson ; Spearman ), suggesting increasing computational overhead as context accumulates. User response times show minimal linear association with turn position on both platforms. These divergent temporal dynamics highlight fundamental architectural differences in how the two systems handle extended interactions.
To evaluate whether verbosity alone provides a reliable predictor of user processing time, we additionally examine correlations between LLM message length and subsequent user response latency. Figure 7 shows that for both Grok and ChatGPT, mean user response time increases with LLM response length, with the steepest growth at shorter outputs and a plateau emerging beyond approximately 4–6k characters. The shaded regions (±1 SD) reveal substantial variability across users. Amongst 6k to 10k the variance further widens and this suggests increased heterogeneity in how users engage with lengthy outputs. This trend suggests that while longer responses are associated with higher average response times in aggregate, length alone is insufficient to predict individual-level latency, with more details in Appendix A.8.
@@ 5 Discussion
The findings presented in this study offer several important implications for understanding human-AI interaction and guiding future research on conversational LLMs. While we only have demonstrated three concrete representative analyses, the dataset’s utility extends well beyond these applications.
@@
Complex Interaction Dynamics in Extended Contexts. ShareChat’s longer conversation contexts, averaging 4.62 turns compared to 2.02 in LMSYS-Chat-1M zheng2023lmsys, provide a critical resource for studying phenomena that emerge only over extended interactions. Prior work has shown that model reliability can degrade as instructions are refined across turns laban2025llms, yet most existing corpora offer limited opportunities to investigate such dynamics due to their predominantly single-turn structure. The presence of conversations extending into dozens of turns, combined with the diversity of intention counts per conversation, enables researchers to examine how user needs evolve, how models maintain coherence over long contexts, and when breakdowns occur. ShareChat makes it possible to study these partial successes systematically, revealing where assistants provide useful but insufficient information and how users adapt their follow-up requests in response. This diversity provides the training signals necessary for models to learn adaptive strategies, specifically, when to shift focus, how to revisit unresolved intentions, and how to maintain coherence as user goals evolve.
@@
Platform-Dependent Usage and Performance Variability. Unlike the “universal assistant” assumption often found in general benchmarks, our cross-platform analysis reveals that user behavior is highly sensitive to the specific design and functional positioning of each platform. Our topic distribution analysis demonstrates that users align their requests with perceived platform strengths, such as using Perplexity for information retrieval versus Claude for technical help. This functional specialization is mirrored in conversation completeness that while ChatGPT and Claude exhibit narrow, concentrated completeness scores, Gemini and Grok show substantial variability. Notably, the high frequency of “partial” verdicts in Perplexity interactions reflects its distinct role as a search engine, where outputs often serve as intermediate steps rather than final answers. Researchers can use ShareChat to investigate these specialized performance profiles, moving beyond generic comparisons to determine which features drive reliability for specific user intents.
@@
Evidence-Based Generation and Temporal Metadata. The preservation of rich metadata enables future research on retrieval-augmented generation (RAG) and interaction pacing in naturalistic settings. Our source content analysis reveals pronounced differences in how platforms integrate external evidence. Grok relies overwhelmingly on concentrated source types, while Perplexity draws from a diversified set of domains. These patterns have implications for understanding how users interpret and rely on cited evidence, and how platform affordances shape trust and verification behaviors. Additionally, the availability of turn-level timestamps enables the investigation of conversational rhythm and interaction pacing, offering a temporal dimension of user experience. Our analysis exposes contrasting architectural dynamics: while ChatGPT demonstrates progressively lower latency as conversations extend over multiple turns, Grok exhibits an inverse trend, with users experiencing slower generation speeds as context length grows.
@@
Self-Selection and Safety in Public Sharing. Finally, the substantially lower toxicity rates in our corpus compared to existing benchmarks, for both user messages and LLM responses, likely reflect the self-selection inherent in public sharing. Users who choose to share conversations may filter for interactions they consider appropriate or valuable, producing a corpus that captures more constructive usage patterns. This suggests that ShareChat offers a cleaner, high-quality resource for studying successful task completion and constructive human-AI collaboration.
@@ 6 Conclusion
In this study, we address the limitations of existing datasets by constructing ShareChat, a large-scale, multi-platform corpus of user-LLM conversations. By compiling more than 140,000 publicly shared interactions from five distinct platforms, we provide a resource that preserves the native metadata and diverse interaction designs of the current AI ecosystem. The dataset’s basic statistics show that the self-selection process inherent in public sharing leads to a corpus with significantly longer conversation contexts and lower toxicity.
We further illustrate the value of this data through three specific analyses. First, our conversation completeness analysis indicate that user intentions are more frequently satisfied in general purpose models like ChatGPT and Claude, reflecting their emphasis on broad capabilities. Second, our source content analysis uncovers the distinct information retrieval strategies employed by Grok and Perplexity. Third, our temporal analysis suggests that ChatGPT and Grok exhibit contrasting temporal dynamics.
Collectively, these findings highlight pronounced differences in user behavior that align with the functional emphasis of each platform, underscoring the importance of studying LLM capabilities within their specific contexts. ShareChat not only provides a testbed for evaluating model reliability across turns, but also offers rich supervision for training models to handle complex, multi-turn user needs in different settings.
@@
Selection Bias. First, the reliance on publicly shared URLs introduces a trade-off between two forms of bias. On one hand, this approach mitigates the Observer Bias which we talked about in Section 1. On the other hand, it introduces Self-Selection Bias heckman1979sample, because users voluntarily choose which conversations to generate links for, the corpus is likely skewed towards interactions they deem successful, interesting, or novel. Therefore, ShareChat may underrepresent mundane utility tasks or failed interactions where the model performed poorly, as users are less motivated to share such content publicly hu2017selection.
@@
Proxy Nature of Automated Metrics. Third, while our conversation completeness analysis and topic analysis utilize robust LLM-based evaluation, it remains a proxy metric. Automated judges provide a scalable approximation of user satisfaction but cannot perfectly capture the nuanced, subjective quality of human preference. These results should be interpreted as strong signals of structural completeness rather than absolute ground-truth measurements of user happiness. Future research is needed to validate these automated metrics against human judgment and to fully explore the evolving longitudinal dynamics of user behavior.
Preliminary Scope of Analysis. Finally, the specific analyses presented in this work, such as the source domain distribution and temporal engagement patterns, are intended as illustrative representative analyses to demonstrate the dataset’s utility. They are not exhaustive sociological studies. Future research is needed to fully explore these dimensions, particularly how these patterns evolve longitudinally as platform features change.
@@
Privacy Preservation and De-Identification. Despite the public provenance of the data, we prioritized the protection of user privacy through a rigorous de-identification pipeline designed to mitigate risks associated with inadvertently exposed sensitive information. First, we utilized automated NER and regular expression matching to detect and redact PII, including real names, email addresses, phone numbers, and IP addresses, for different languages. In addition, we do not store or release original platform-specific user IDs or usernames. All user identifiers were cryptographically hashed and converted into internal project identifiers during the initial processing stage. Finally, the analyses presented in this paper, like the toxicity assessments and completeness scoring, are conducted solely on aggregated statistics. No individual user’s conversation history is singled out for qualitative scrutiny in a manner that could lead to re-identification or personal embarrassment.